the supplement of an angle measures 25 degrees more than twice it's complements. find the measure of the angle
Answer by nerdybill(7384)

the supplement of an angle measures 25 degrees more than twice it's complements. find the measure of the angle
.
Let x = measure of the angle
then
90-x = complement
180-x = supplement
.
180-x = 2(90-x) + 25
180-x = 180-2x + 25
-x = -2x + 25
x = 25 degrees